Experiment Startup activities Both the machine as well as the Brahms setup has come to a stage where we can take some useful date to debug and setup the detector. The local BNL and Bucharest people will carry out most, if not all of this early work. This afternoon we discussed what kind of progress could be achieved over the weekend, and what problems identified before the Monday access. There was a short store this afternoon with some data recorded. Known issues The FFS1440 and Bfs1440 cannot be talked to by Brahvo. Thus H1,H2,C1, new FS trigger counters cannot be turned on. TPM2 has a gating pulser connected. This is because the mtp2 readout board was faulty. Today Bob identified the problem and fixed it (a faulty switch on the receiver board). TPM2 can now be read out needs the pulser cal done. The earliest time for doing this is Monday. The ZDC looks reasonable (somehow (Michael!?) had interchanged the left and right signal in the FEH so the BB and ZDC vertex pos was anti-correlated). The changed HV-module for tpc's in L1458A have a badly calibrated resistor so the drift field is restricted to 4450V (should be 4950). Will be fixed on Monday. Check today. MTP1 looks reasonable- the event size for ZDC events were rather large, back to ~ 20K per sequence. This obviously have implication for triggering and dead time. The readout time for MTP1 is ~ 8 msec. TOFW looks quite good, as does BB and ZDC TFW-2 HV setup has to be modified. The delay cabling for the new trigger counters were done today (John, Catalin, Bob). These are ready to include in the data stream. Over this weekend will address the following experiment setup issueswith main emphasis to identify problems that can be fixed/addressd on the Mondat=y access. a.. Pedestal T1 (that seems Off for lower tie bins (20-40)) b.. Turning on T2 , pedestals + data c.. Tiles setup + Si (Hiro) d.. Analyze data-runs to calibrate offsets BB,ZDC. e.. TDC time calibrations. f.. First settings of DC T3,T4, and T5 g.. Systematic zero field runs for FFS a.. D1=0, D2=1/8,1/4 b.. D1=0,D2-0 c.. D1=1/8,1.4, D2=0 d.. Both on. Results will have to be looked upon. Since the machine performance is still quite questionable the plan is roughly a.. Friday night: Hiro (emphasis on Tile+Si) b.. Saturday night: T2, possibly zero files running c.. Sunday: some systematic longer runs to get data for looking more carefully on data. Non-beam related Update readout map Update HV files.. Update Supermon, Ntuples. Thanks to Hiro, JH, Catalin, Oana who will be available for these activities. Flemming ---------------------------------------------------------------- Flemming Videbaek Physics Department Brookhaven National Laboratory e-mail: videbaek@bnl.gov phone: 631-344-4106 _______________________________________________ Brahms-l mailing list Brahms-l@lists.bnl.gov http://lists.bnl.gov/mailman/listinfo/brahms-lReceived on Fri Dec 19 19:34:47 2003
This archive was generated by hypermail 2.1.8 : Fri Dec 19 2003 - 19:35:05 EST